Night report 2021-08-05

Observers: Matt Nicholl, Evan Ridley, Ben Gompertz

Support: Sara Muñoz, María Delgado, Raúl Gonzalez, Lluís Galbany

TAT: Jacob Teffs, Michael Fulton, Janet Chen

EFOSC Calibrations:

Afternoon: pessto_calib

(Morning: pessto_bias)

Weather:

Start: humidity ~ 15%, wind ~ 4m/s, seeing ~ 1.1"

Middle: humidity ~ 11%, wind ~ 8m/s, seeing ~ 1.4"

End: humidity ~ 11%, wind ~ 12m/s, seeing ~ 1.4"

Comments (times in UT):

Image analysis done after first standard

05:50 - LOEM failed due to network problem at La Silla, persevering over teams - target acquisition via TO's phone!!